OpenStudy (anonymous):

the area of a circular rose garden is 22 square meters. what is the radius of the garden? o_O thats all it gives me i dont get it

OpenStudy (campbell_st):

well the formula for the area of a circle is \[A = \pi r^2\] and you know A = 22 and pi is a number so \[22 = \pi r^2\] the steps are... 1. divide both sides of the equation by pi 2. take the square root to find r.

OpenStudy (anonymous):

\[Area _{circle}=\pi r ^{2}\] What you know; \[Area _{circle}=22m ^{2}\] \[\pi \approx3.14\] What you want to know; \[r=?\] Process; Plug in values. \[22=3.14r ^{2}\] Now you need to use the Order of Operations in reverse to get r alone; \[\frac{ 22 }{ 3.14 }=\frac{ 3.14r ^{2} }{ 3.14 }\] \[22\div3.14\approx7\] \[\sqrt{7}=\sqrt{r ^{2}}\] \[\sqrt{7}\approx2.65\] \[2.65 = r\] \[r=2.65m\]